Circular No. 7793

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002C IN IC 3376 B. Swift and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Jan. 14.5 (mag 16.7) and 15.5 UT (mag 16.8). SN 2002C is located at α = 12^{h}27^{m}52^{s}.66, δ = +27^{o}00′06′′.1 (equinox 2000.0), which is 30′′.6 east and 30′′.7 north of the nucleus of IC 3376. A KAIT image taken on 2001 May 27.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001ig IN NGC 7424 A. Clocchiatti, Pontificia Universidad Católica de Chile, reports on

his spectroscopic observations of SN 2001ig obtained on Jan. 13.0 UT with the European Southern Observatory 3.6-m telescope (+ dual imager/spectrograph EFOSC2): “A fully reduced spectrum (336–1030 nm; resolution 0.9 nm) with good signal-to-noise shows that SN 2001ig is, as expected (cf. Phillips et al., IAUC 7772; Clocchiatti and Prieto, IAUC 7781), making a transition from type II initially to type Ib/c now. The spectrum resembles that of SN 1993J some 40 days after shock breakout (Swartz et al. 1993, Nature 365, 232), although the spectrum of SN 2001ig displays stronger Ca II lines, weaker H and He I lines, and faster expansion velocities. The Hα P-Cyg line shows the characteristic complex shape of the transitional stage (Filippenko et al. 1994, A.J. 108, 2220). As expected, the blueshifted minimum is blended with the Si II 635.5-nm line and the redshifted maximum indented by the P-Cyg profile of the He I 667.8-nm line. The positions of blueshifted minima, proposed line identifications, and resulting approximate expansion velocities (assuming a recession velocity of 942 km/s for NGC 7424, from the NASA Extragalactic Database) are: 380 nm, Ca II (H and K), 12 000 km/s; 469 nm, Hβ, 11 900 km/s; 478 nm, He I (492.1 nm), 9800 km/s; 487 nm, He I (501.5 nm), 10 000 km/s; 570 nm, He I (587.6 nm, blended with Na I D), 10 000 km/s; 628 nm, Hα (blended with Si II 635.5-nm), 14 100 km/s; 647 nm, He I (667.8 nm), 10 600 km/s; 830 nm, Ca II (849.8, 854.2, and 866.2 nm), 22 500 km/s. The O I 777.4nm line is also detected, with a blueshifted minimum near 760 nm. The center of the profile cannot be determined, since it is severely blended with a strong telluric absorption band. Observations of this strange supernova are urged, since its solar elongation is now getting very small.”

2002 January 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green